Looking atthis pair of linear equation, what would your firs stp? (Write your answer and model it )
1. 2y = 6x - 4
-2 + y = 3x

Need to graph it after

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

Looking at this pair of linear equation, what would your first step? (Write your answer and model it )
1.
2y = 6x - 4
-2 + y = 3x
---
1st step: Solve each equation for "y":
y = 3x-2
y = 3x+2
----
2nd step: Plot 2 points for each equation and draw the two lines.
3rd step: Find the point of intersection.
Comment: Your two lines do not intersect; they are parallel.
==============================================================
Cheers,
Stan H.




2. 27x + 9y = -18
-12x + 4y = 8
Need to graph it after

Answer by richwmiller(17219)

2y = 6x - 4
-2 + y = 3x
6x-2y=4
3x-y=-2
First step is to put them in the same order. I chose standard x then y format.
Find slopes
-a/b=-6/-2=3
-a/b=-3/-1=3
So they are parallel and do not have a solution.